What mass of C4H6 must react to produce 43.2 L of CO2 gas at 499 K and 2.57 atm. 

2C4H6(g)+11O2(g)⟶8CO2(g)+6H2O(l)

 

First, use the ideal gas equation to calculate the moles of CO2.

Then, calculate the mass of C4H6 from the moles of CO2. Show the conversions required to solve this problem.


1
Expert's answer
2022-04-14T07:20:02-0400

V = 43.2L, T = 499K, P = 2.57atm

R = 0.0821 L*atm/moll*K

PV = nRT

n = PV/RT, n(CO2) = (2.57*43.2) / (0.0821*499) = 111.024/40.9679 = 2.71 moll


m(C4H6) = (2Mr(C4H6)*V(CO2)) / 8Vm(CO2)

V(CO2) = n(CO2) * Vm(CO2), V(CO2) = 2.71*22.4 = 60,7 L

m(C4H6) = ((2*54)*60.7)/8*22.4 = 36.58 g